Given that 2−3i is a root of the equation z^3+pz^2+qz+13 = 0, find the other two roots and the values of the real constants p and q.
1
Expert's answer
2020-05-20T16:47:05-0400

"(a+b)^3=a^3+3a^2b+3ab^2+b^3=>(2-3i)^3=8-36i-54+27i=-46-9i\newline
(2-3i)^2=4-12i-9=-5-12i"

So we have:

"-46-9i+p(-5-12i)+q(2-3i)+13=0\newline
(2q-5p)+(-12p-3q)i=33-9i=>\newline
\begin{cases}
2q-5p=33\\
12p+3q=-9
\end{cases}\newline


\begin{cases}
2q-5p=33\\
q=-3-4p
\end{cases}\newline


\begin{cases}
-6-8p-5p=33\\
q=-3-4p
\end{cases}\newline

\begin{cases}
p=-3\\
q=9\\
\end{cases}\newline"

It means that:

"z^3-3z^2+9z+13=0"

"\dfrac{z^3-3z^2+9z+13}{z-2+3i}=z^2+(-1-3i)z+(-2-3i)"

"z^2+(-1-3i)z+(-2-3i)=0\newline
D=(-1-3i)^2-4(-2-3i)=1+6i-9+8+12i=18i\newline
\sqrt{D}=3\sqrt{2i}\newline
if \sqrt{i}=\dfrac{1}{\sqrt{2}}+\dfrac{1}{\sqrt{2}}i, thenD=3+3i\newline
if \sqrt{i}=-\dfrac{1}{\sqrt{2}}-\dfrac{1}{\sqrt{2}}i, thenD=-3-3i\newline
D=3+3i\newline
z_2=\dfrac{1+3i+3+3i}{-2-3i}=-2\newline
z_3=\dfrac{1+3i-3-3i}{-2-3i}=\dfrac{2}{2+3i}\newline
D=-3-3i\newline
z_2=\dfrac{1+3i-3-3i}{-2-3i}=\dfrac{2}{2+3i}\newline
z_3=\dfrac{1+3i+3+3i}{-2-3i}=-2\newline"